OpenStudy (anonymous):

1+2b+12=5b-5 explain your process please

OpenStudy (anonymous):

1+2b+12=5b-5 add the 5 to the other side and you have 6+2b+12=5b then u subtract 2b by -2b and add it to the other side like 6+12=3b and then add 6+12 to get 18 and divide 18 by 3 to get 6.